ln(5x-2)=12. approximate the solution to four decimal places.
Answer by edjones(7569)

ln(5x-2)=12
e^(ln(5x-2)=e^12
5x-2=162754.791419
5x=162756.791419
x=162756.791419/5
x=32551.3583
.
Ed